OpenStudy (anonymous):

A rectangle park measures 300 ft by 400 ft. A sidewalk runs diagonally from one corner to the opposite corner. Find the length of the sidewalk. a. 400 ft c. 250 ft b. 500 ft d. 650 ft

OpenStudy (mathmale):

Since the park is rectangular, it has 4 90-degree interior angles. Connecting any corner to the diagonally opposite corner is the same as drawing the HYPOTENUSE of a right triangle inside the park boundaries. Does that help you visualize the layout? Draw a picture and label whatever it is that you know.

OpenStudy (zpupster):

 from pythagorean theorem a^2 + b^2= c^2 there is a famous triangle 3-4-5 3^2 + 4^2=5^2 9+16=25 what if i multiply all these values by 100??

OpenStudy (mathmale):

zpupster is one up on me. Nice work! Now, Nathan, what next?

OpenStudy (anonymous):

I need to find the hypotenuse hold I now how to do that

OpenStudy (mathmale):

You're right on target. Please review the Pythagorean Theorem.

OpenStudy (anonymous):

Ok so I got 250000 now I need to find the square root of it and I'll have my answer. Ugh big numbers give me headaches

OpenStudy (anonymous):

It's b 500! Right?

OpenStudy (zpupster):

500*500=250000

OpenStudy (zpupster):

yes!!

OpenStudy (anonymous):

Alright a! Thank you!

OpenStudy (zpupster):

look for the 3-4-5 in a lot prolem or multiples of 6-8-10 9-12-16 12-16-20 etc. etc. etc.